OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates ordered pair negative 3 comma 5 pi divided by 3 for my answer I keep getting -3, -1

OpenStudy (anonymous):

but these are my choices ordered pair negative 3 divided by 2 comma 3 divided by 2 ordered pair 3 square root 3 divided by 2 comma negative 3 divided by 2 ordered pair negative 3 divided by 2 comma 3 square root 3 divided by 2 ordered pair 3 divided by 2 comma negative 3 divided by 2

OpenStudy (mathstudent55):

\((r, \theta ) = (x, y) \) \( x = r \cos \theta \) \( y = r \sin \theta \) Since your \(\theta\) is \( \dfrac{5 \pi}{3} \), set your calculator to radians before taking the sine and cosine.

OpenStudy (anonymous):

si is it this ordered pair 3 square root 3 divided by 2 comma negative 3 divided by 2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

\(\left(3 \cos \dfrac{5 \pi}{3}, 3 \sin{\dfrac{5 \pi}{3}} \right)\) \( \left( 3 \times \dfrac{1}{2} , 3 \times \left(\dfrac{-\sqrt{3}}{2}\right) \right) \) \( \left( \dfrac{3}{2} , - \dfrac{3\sqrt{3}}{2}\right) \)

OpenStudy (ranga):

If r = -3 in the problem, the answer above with signs switched: x negative and y positive.

OpenStudy (mathstudent55):

@ranga You are correct, the 3 in the problem is actually -3, so just take the negative of the x- and y-coordinates I calculated. @llkramer In the future, write numbers as numbers. Don't write "negative three." Instead, write "-3." It's much easier to read.

OpenStudy (anonymous):

quick question I get this now thank you :) but for some reason when I logged on open study it doesn't allow me too see the equation input system. you guys know anything about that?

OpenStudy (mathstudent55):

The equation editor shows on the left bottom of the reply area.

OpenStudy (anonymous):

I know I can use it but when I post them or someone else post them its in form like this sqrt5/pie2lim2^2

OpenStudy (mathstudent55):

Then your browser is not displaying equations properly. That has happened to me. Somtimes by closing open study and starting it again, it displays the equations properly again.

OpenStudy (anonymous):

You're cute